OpenStudy (cartoon_horses):

what are the center and radius of the circle defined by the equation x2+y2-6x+10y+25=0? I am halfway done with this and I'm stumped by this question

OpenStudy (radar):

Review this link, and you will be on your way. http://www.purplemath.com/modules/sqrcircle.htm

OpenStudy (sshayer):

can you show me your work?

OpenStudy (cartoon_horses):

Honestly, I just search up half of my answers, i rarely ever work out the problems

OpenStudy (sshayer):

\[\left( x^2-6x+? \right)+\left( y^2+10y+? \right)=-25+?+? \]

OpenStudy (sshayer):

complete the squares \[\left( \frac{ -6 }{ 2 } \right)^2=9\] \[\left( \frac{ 10 }{ 2 } \right)^2=25\]

OpenStudy (cartoon_horses):

These are the answer choices, Center 3,-5 Radius 9 Center -3,-5 radius 3 Center 3,-5 radius 3 Center -3,5 radius 9

OpenStudy (sshayer):

\[\left( x^2-6x+9 \right)+\left( y^2+10y+25 \right)=-25+9+25\] \[\left( x-3 \right)^2+\left( y+5 \right)^2=3^2\] which is of the form \[\left( x-h \right)^2+\left( y-k \right)^2=r^2\] center is (h,k) and radius r

OpenStudy (sshayer):

can you find the correct option ?

OpenStudy (cartoon_horses):

Yes

OpenStudy (cartoon_horses):

Thank you so much

OpenStudy (sshayer):

yw